OpenStudy (anonymous):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (nikato):

any ideas?

OpenStudy (anonymous):

no it really confuses me!!

OpenStudy (nikato):

ok. lets look at your problem. it says equation 2 multiplied by 1. which answer choice has equation 2 multiplied by 1?

OpenStudy (anonymous):

isn't it the second choice?

OpenStudy (nikato):

no. the second equation is 2x-3y=1 and if you multiply 1, you'll still get the same equation. and the only choices that has this equation is A and D

OpenStudy (anonymous):

no never mind... I mean the last one because the 2nd equation stayed the same cause its multiplied by 1

OpenStudy (nikato):

now you have to decide btw A and D the problem says " to replace equation 1 with the sum of equation 1 and 2" so you wanna add the 2 equations

OpenStudy (nikato):

 and what do you get?